In Denmark, a survey of 352 individuals found that 131 were homozygous for Met, 168 were heterozygous, and 53 were homozygous for Val. Is the PRNP gene in Hardy-Weinberg equilibrium? Show your calculations. (4 marks)
1
Expert's answer
2016-10-15T09:16:03-0400
The frequency of allele for Met is (2*131+168)/2*352=0.61
The frequency of allele for Val is (2*53+168)/2*352=0.39
According to the calculated frequencies, the number of individuals, who are homozygous for Met, should be: 0.61^2 *352=131
In the same way, the expected number of heterozygous individuals is 2*0.61*0.39*352=168
Similarly, the expected number of individuals who are homozygous for Val is: 0.39^2*352=53
Thus, the PRNP gene is in Hardy-Weinberg equilibrium